you are considering investing in a security that will you $5100 in 32 years. assume these investments is worth $654 today; what is the rate of return investors earn on this investment
Answer #1

Calculate the rate of return on investors as follows:

* Je=RATE(32,-654,5100) Function Arguments ? X RATE Nper 32 Pmt Pv -654 Fv 5100 Type = 32 = number = -654 = 5100 * = number =

Therefore, the rate of return is 6.63%.

------------------------------------------------------------------------------

Alternative approach:

Future value = Present value*(1+rate)^years

$5100 = $654 *(1+ rate)^32

(1+rate)^32 = $5100/$654

Rate = ((7.79816514)^(1/32)) -1

Rate = 1.0663 - 1

Rate = 0.0663 or 6.63%.